OpenStudy (anonymous):

M(4, 2) is the is the midpoint of RS. The coordinates of S are (6, 1). What are the coordinates of R?

sam (.sam.):

Do you remember the midpoint formula?

sam (.sam.):

\[\huge M(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}) \\ \\ \huge M(\frac{x_1+6}{2}=4,\frac{y_1+1}{2}=2) \\ \\ \huge \frac{x+6}{2}=4~~~~~~and~~~~~~\frac{y+1}{2}=2\] Solve